OpenStudy (anonymous):

solve: cosx+1=sinx, [0,2π)

OpenStudy (callisto):

\[\cos x+1=\sin x\]\[\cos x - \sin x = -1\]By squaring both sides, what do you get?

OpenStudy (zzr0ck3r):

clever:)

OpenStudy (callisto):

I'm too stupid to think of other tricks :(

OpenStudy (zzr0ck3r):

lol:)

OpenStudy (jhannybean):

There is nothing stupid abut that. It's the easiest method.

hartnn (hartnn):

yup, nice way.... other longer method would be to square both sides initially to get a quadratic in cos^2 x....

OpenStudy (zzr0ck3r):

or just look at it:)

OpenStudy (anonymous):

... I think I'm really missing it right now because I don't get what seems so easy to you T_T... Haha sorry but if you squared it like so: (cosx-sinx)^2=1, wouldn't it become more complicated? cosx^2-2cosxsinx+sinx^2=1 What do I do then?

OpenStudy (jhannybean):

\[(\cos x - \sin x)^2 = (-1)^2\]

hartnn (hartnn):

use the most standard identity of trigonometry \(\large \sin^2x+\cos^2x=1\)

OpenStudy (callisto):

Something you need: \[cos^2x+sin^2x =1\]\[2sinxcosx=sin(2x)\]

OpenStudy (jhannybean):

\[(\cos x - \sin x)^2 = (-1)^2\]\[(\cos x-\sin x)(\cos x - \sin x)=(-1)(-1)\]can you FOIL and simplify?

OpenStudy (anonymous):

yeah I did, I understand now. Thanks :)

OpenStudy (anonymous):

wait, one last question though. Is there a difference between \[\cos ^{2}x\] and \[cosx ^{2}\]

OpenStudy (callisto):

Yes.

OpenStudy (callisto):

\[cos^2x = (cosx)^2\]You take the cosine before taking the square. \[cosx^2=cos(x^2)\]You take the square before taking the cosine.

OpenStudy (anonymous):

Thank you! :) I was planning to use that identity too but I thought that if you squared both side it would become cos(x^2) and that didn't match the identity. :O But thanks for clearing it up!!

OpenStudy (callisto):

Welcome~~~~~ :)